OpenStudy (anonymous):

A spinner contains 4 equal sections, one of which is yellow, and the rest are blue. If the spinner is spun 3 times, determine the probability that: a) it lands on yellow exactly twice b) it lands on yellow at least twice

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

OKay, on this I'm trying to figure out, the probability all together is 4/6 right? So simplify that and see what you get!

OpenStudy (zarkon):

do you know the prob of getting yellow?

OpenStudy (zarkon):

on one spin

OpenStudy (anonymous):

no i'm not sure

OpenStudy (anonymous):

IN your question it says there's 4 sections, 1 yellow the rest blue, which makes 3 blues, (SOrry I told you differntly earlier)

OpenStudy (anonymous):

Do you get that?

OpenStudy (anonymous):

ya all are 4 but how u got 6

OpenStudy (anonymous):

Sorry, that was wrong!!

OpenStudy (anonymous):

I typed the wrong number!

OpenStudy (anonymous):

what i did is 2/4 = 1/2 but i'm not sure for my answer

OpenStudy (anonymous):

cuz it exactly twice yellow in a

OpenStudy (anonymous):

is that true

OpenStudy (anonymous):

I" mnot sure, trying to figure it out!

OpenStudy (anonymous):

Sorry, I thought I had it, but had to re look

OpenStudy (anonymous):

yes please

OpenStudy (zarkon):

I would start by answering my question

OpenStudy (anonymous):

i need ur help

OpenStudy (anonymous):

the prob of getting yellow is that 2/4 i'm not sure

OpenStudy (zarkon):

# of yellow sections/total number of sections

OpenStudy (anonymous):

oh it means 1/4

OpenStudy (anonymous):

right

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

what is the prob you don't get yellow then

OpenStudy (anonymous):

i didn't get it

OpenStudy (zarkon):

ie what is the probability of getting a blue

OpenStudy (zarkon):

you will need both of these numbers

OpenStudy (anonymous):

3/4

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

so if you want exactly 2 yellow that can happen 3 different ways ...

OpenStudy (zarkon):

YYB YBY BYY

OpenStudy (zarkon):

so you need to compute the prob of getting YYB in that order

OpenStudy (anonymous):

there is only 1YYB so it means 1/3

OpenStudy (zarkon):

no

OpenStudy (zarkon):

if each spin is independent then you need to multiply the probabilities of getting Y and Y and B

OpenStudy (zarkon):

ie \[\frac{1}{4}\cdot\frac{1}{4}\cdot\frac{3}{4}\]

OpenStudy (zarkon):

are you studying the binomial distribution?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

solving prob problem

OpenStudy (zarkon):

ok...this problem follows a binomial distribution...just so you know

OpenStudy (zarkon):

so YYB gives \(\displaystyle\frac{3}{64}\)

OpenStudy (zarkon):

I assume you have studied independent events.

OpenStudy (anonymous):

this is for a right what about b

OpenStudy (zarkon):

we are not done with (a)

OpenStudy (anonymous):

oh

OpenStudy (zarkon):

you still need to do YBY and BYY

OpenStudy (anonymous):

it will be 3/64

OpenStudy (zarkon):

yes...for each

OpenStudy (zarkon):

so what is the prob of getting exactly 2 Yellow?

OpenStudy (anonymous):

we add all of it 3/64 + 3/64 + 3/ 64

OpenStudy (zarkon):

correct and get 9/64

OpenStudy (zarkon):

now we can do part (b)

OpenStudy (zarkon):

at least twice means 2 or more times

OpenStudy (anonymous):

thnx ya

OpenStudy (zarkon):

so we know the prob of exacty 2 ...what is the prob of exactly 3 yellow?

OpenStudy (anonymous):

we add on it the same #

OpenStudy (zarkon):

it will not be 3/64

OpenStudy (anonymous):

YYY

OpenStudy (anonymous):

like this

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

so 3/729

OpenStudy (zarkon):

\[\frac{1}{4}\cdot\frac{1}{4}\cdot\frac{1}{4}\]

OpenStudy (anonymous):

1/64

OpenStudy (zarkon):

yes...now add to our previous result

OpenStudy (anonymous):

9/64 + 1/64= 10/ 64

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

oh thnx I really appreciate that's so helpful

OpenStudy (zarkon):

np

OpenStudy (anonymous):

can i ask u 1 more quetion i need some one 2 understand me like that